Show all work. Label and clearly explain your answer. This is very important. 1) You must explain how you arrived at your answer in order to get full credit. 2) If you do show your work, and your answer is wrong, you can still earn a substantial amount of credit depending on how serious the error is. 3) If your answer is wrong, and you don't show your work, you will get a zero. Sarcosuchus Corporation offers very short-term loans to applicants with low credit ratings. They charge an interest rate of 0.03333 percent per hour. a. If you borrow $100 from Sarcosuchus, how much will you owe in one week? b. If you borrow $100 from Sarcosuchus, how much will you owe in one month (30 days)? c. What annual rate is Sarcosuchus charging? Hint: Convert the simple hourly rate of 0.03333 percent into simple annual rate
Answer #1

B 1 Principal Period rate Interest Amount owe in one week 100 168 0.03333% 5.5994 105.5994 10 11 Principal Period rate Intere

E17 - Α. в 60 vu WNA Principal Period rate Interest Amount owe in one week 100 =24*7 0.0003333 =C4*C6*C5 =SUM(C4,C7) 11 12 Pr